OpenStudy (nick63689):

On average, a commercial airplane burns about 3.9 × 10³ mL of gasoline per second. There can be as many as 5.1 × 10³ airplanes in the air in the United States at any given time. About how many milliliters of gasoline are burned per minute by commercial airplanes in flight in the United States? 1.2 x 10^( )mL So 1.2 x 10 to what power?

OpenStudy (nick63689):

@hartnn @TheSmartOne @Jadeishere @welshfella @Nnesha @JoannaBlackwelder @ganeshie8 @JFraser @Loser66 @Vuriffy @skullpatrol @Apple @ILovePuppiesLol @imqwerty @AloneS @Ambuj @agent0smith @jhonyy9 @mathstudent55 @whpalmer4 @Wolf @Ellie201 @rayep12 @rosielopez_17 @Revircs @rbrasch @Tifferslovesyaoi @TinaSh @Ttrichards @todjunior @ttttttt @YanaSidlinskiy @yasatg @UnkleRhaukus @Unavailabilityy @userk2 @oliviad @oopsydaisy35 @poopsiedoodle @PandaNinja101 @procrastilate @pooja195 @Agl202 @Aleah54 @AAbomosalam1998 @Autumn_12001 @Ashton_Bishop @anastasiai @sweetburger @samanthagreer @something @sam_OVO @smollberru @sophiesky @Daniellelovee @Dyann.Cardona @ddoll01 @Deadpool @Finette @gudetamas @Goldendragon01 @Hayhayz @HocusPucus292333393 @hadeel @htxatlarge @jtug6 @johnsmith31613 @jjjjjjjjjjjjhujljnjb @Jaynator495 @jim_thompson5910 @jabez177 @kewlgeek555 @kidrah69 @kellyspeakslouder @k_d_1 @kitten_dude @leenathan @Lacris @lyss @ljetibo @Lyssburleson @Loli @zzr0ck3r @zepdrix @basically @breanna.rae @Batman4ever @nerdkid @noma @MarcoReus11 @mantar0078 @Makayla_Caldwell @kevin @will.h

OpenStudy (nick63689):

Le mass tagging .3.

OpenStudy (yanasidlinskiy):

Like the whole US to the rescue?

OpenStudy (nick63689):

mhm

OpenStudy (leenathan):

IM BEING TAGGED TO MUCH

OpenStudy (yanasidlinskiy):

Twice as many commercial aircraft burn twice as much gasoline, so we need to multiply how much 1 airplane uses by the amount of airplanes in the sky. In this case, the problem tells us this is 5.0 x 10^3. (3.9 x 10^3) x (5.1 * 10^3) = 20.0 * 10^6, which equals 2.0 * 10^7 mL/sec. ^^^ This answer is rounded... We may appear to be done here, but we need to be careful; the problem is asking us how much is burnt every MINUTE, not every second. A minute contains 60 seconds, so that means 60 times as much gasoline is burnt. 60 x (2.0 * 10^7) = 120.0 * 10^7, which equals 1.2 * 10^9 by scientific notation. Thus our answer is 1.2 * 10^9 mL/min of gasoline.

OpenStudy (yanasidlinskiy):

Lee, now imagine Ambassadors with all the mass tags .-. Why should you complain?

OpenStudy (nick63689):

THANK YE! Heres a joke to show my appreciation: Why did the Knight stop using the internet? He was tired of chainmail. *silence*

OpenStudy (yanasidlinskiy):

By the way Nick, you might want to get second answers. I'm actually doing my Anatomy and kinda did math in my book lol but umm yea.. definitively get it checked. LOL

OpenStudy (nick63689):

I just needed help with that one question, and you got it right xD

OpenStudy (yanasidlinskiy):

Oh dang hahha. Phew .-. Sorry Anatomy is just killing me to the core.

OpenStudy (nick63689):

MUSIC is killing me to the core .o.

OpenStudy (yanasidlinskiy):

Ha, anatomy is worst. Only if you knew the struggle... Smh... Like I'll do anything to get away from Anatomy right now

OpenStudy (nick63689):

"Which best describes Idiophone" Me: wat.

OpenStudy (yanasidlinskiy):

\Bleh.

OpenStudy (nick63689):

Bleeehhh

OpenStudy (yanasidlinskiy):

Heck yea

OpenStudy (zzr0ck3r):

super annoying

OpenStudy (zzr0ck3r):

Can we have these users that tag the whole site removed? I hate narcissist.

OpenStudy (leenathan):

k yana next time tag me XD

OpenStudy (zzr0ck3r):

the questions are listed on the left side of your screen. If you can help....then do it.